OpenStudy (anonymous):

Which ordered pairs are solutions to the equation? y = 3x – 4 A. (11, 5) and (2, 2) B. (5, 11) and (3, 2) C. (5, 11) and (2, 3) D. (5, 11) and (2, 2)

OpenStudy (anonymous):

@HazelLuv99

OpenStudy (texaschic101):

time to do some subbing... y = 3x - 4 I will try the first answer choice...(11,5) and (2,2) (11,5)...x = 11 and y = 5 5 = 3(11) - 4 5 = 33 - 4 5 = 29 (incorrect...we do not even need to try (2,2) because the first one (11,5) is incorrect. now you can try the others and see what fits...what is equal...give it a shot :)